1. The largest moon of Saturn, Titian has a mean orbital radius of 9 1.22×10 m . The orbital period of Titan is 15.95 days. Hyperion, another moon of Saturn, orbits at a mean radius of 9 1.48×10 m . Use Kepler’s third law of planetary motion to predict the orbital period of Hyperion in hours.


1
Expert's answer
2021-06-28T18:07:02-0400

According to Kepler's third law, "\dfrac{T^2}{a^3} = \dfrac{4\pi^2}{GM}" . Therefore, for two satellites of the same planet we'll obtain

"\dfrac{T_1^2}{a_1^3} =\dfrac{T_2^2}{a_2^3}" or "\left(\dfrac{T_2}{T_1}\right)^2 = \left(\dfrac{a_2}{a_1}\right)^3" . In such form we may write periods in arbitrary units, so in days

"\left(\dfrac{T_2}{15.95}\right)^2 = \left(\dfrac{1.48\cdot10^9}{1.22\cdot10^9}\right)^3 , \\
T_2 = 1.336T_1 =21.31^d = 511.5^h"